# Psychiatry-Bench: A Multi-Task Benchmark for LLMs in Psychiatry

### v1

@@ 6 Results & Discussion
This section presents the results of our experimental evaluation across multiple psychiatric QA tasks, using a range of LLMs assessed by expert-generated reference answers and LLM-as-a-judge scoring. We analyze performance trends across models and task types, comparing output quality under different prompt strategies and exploring the strengths and weaknesses of each model in tasks such as diagnosis, treatment planning, and follow-up care. Our discussion addresses both quantitative outcomes and qualitative patterns in reasoning, highlighting areas where LLMs demonstrate promising clinical utility and where challenges remain in achieving safe, interpretable, and consistent psychiatric decision support.
6.1 Diagnosis
Table 5: Diagnosis Results: Average Similarity Scores for Model-Generated Diagnosis Outputs.The underlined bold values denote the best score.
Model Name Avg. Similarity score
Deepseek-R1 $84.2$
Deepseek-chat $82.7$
Gemini 2.5 Pro Preview (03-25) 86.0
Gemini 2.0 Flash $83.1$
Llama 3.3 70B $83.6$
QWQ-32 $85.9$
Gemini 2.5 Flash Preview (04-17, Thinking) $85.0$
Gemini 2.5 Flash Preview (04-17) $78.0$
We generated model outputs using the Engineering prompt template 1 and the Evaluation prompt template 2 separately.As shown in Table 6, the average similarity scores, evaluated by a Llama 3.3 70B judge, are reported for eight models on the diagnosis task. The scores range from (78.0%) to (86.0%), indicating that all evaluated models demonstrate strong semantic alignment under the given prompts.The top-performing model, Gemini 2.5 Pro Preview (03-25), achieved the highest similarity score at (86.0%), highlighting the impact of recent architectural refinements or expanded training data. Close behind is QWQ-32, scoring (85.9%), effectively tying for second place and proving competitive despite limited adoption. Gemini 2.5 Flash Preview (04-17, Thinking) followed with a solid (85.0%), showing how variations in inference behavior (e.g., "Thinking" mode) can yield measurable gains in semantic alignment. At the other end of the spectrum, Gemini 2.5 Flash Preview (04-17), without the "Thinking" mode, recorded the lowest score at (78.0%), underperforming relative to its sibling variants. Deepseek-chat scored (82.7%), placing it above the lowest performer but still trailing the top contenders. Mid-tier results included DeepSeek-R1 (84.2%), Llama 3.3 70B (83.6%), and Gemini 2.0 Flash (83.1%), all delivering solid, if not leading, performance. Overall, the results span a compact range of 8.0 percentage points, from (78.0%) to (86.0%), indicating a generally strong level of performance across all models. However, the standout scores from newer or enhanced models, especially Gemini 2.5 Pro Preview, underscore the benefits of continual optimization in both prompt design and model development.
6.2 Treatment
Table 6: Treatment Results: Average Similarity Scores for Model-Generated Treatment Outputs. The underlined bold values are the best score.
Model Name Avg. Similarity score
Deepseek-R1 79.9
Deepseek-chat 81.9
Gemini 2.5 Pro Preview (03-25) 80.5
Gemini 2.0 Flash 77.4
Llama 3.3 70B 82.3
QWQ-32 82.1
Gemini 2.5 Flash Preview (04-17, Thinking) 82.9
Gemini 2.5 Flash Preview (04-17) 81.0
We generated model outputs using the Engineering prompt template 17 and the Evaluation prompt template 18 separately.As show in table  5 all eight models deliver moderately high similarity scores for their treatment recommendations, ranging from (77.4 %) to (82.9 %). The Gemini 2.5 Flash Preview (04-17, Thinking) led the set at (82.9 %), followed closely by Llama 3.3 70B at (82.3 %) and QWQ-32 at (82.1 %). Deepseek-chat achieved (81.91 %), while the secondary Gemini 2.5 Flash Preview (04-17) variant score (81.0 %). Gemini 2.5 Pro Preview (03-25) and Deepseek-R1 recorded (80.5 %) and (79.9%), respectively, with the original Gemini 2.0 Flash trailing at (77.4 %). The spread of (5.49 %) between the top and bottom performers highlights a meaningful performance gap in generating treatment plans. The clear advantage of the Gemini 2.5 Flash Preview (04-17, Thinking) suggests that its latest reasoning capabilities particularly enhance clinical coherence and relevance. Llama 3.3 70B and QWQ-32 also demonstrate that robust language understanding and contextual breadth translate into more accurate treatment outputs, marking them as strong baselines for further development.At the lower end, Gemini 2.0 Flash’s (77.4% ) score underscores how quickly performance can lag without iterative updates, while Deepseek-R1’s sub-80% result indicates potential weaknesses in domain-specific knowledge or recommendation specificity. Deepseek-chat’s solid upper-mid placement (81.9 %) shows that conversational optimization aids treatment advice, but it still falls short of models incorporating the newest architectural enhancements. Overall, these findings point to the importance of continuous model refinement to achieve clinically reliable treatment generation.
6.3 Treatment Follow-Up
Table 7: Follow up Results: Average Similarity Scores for Model-Generated Treatment Follow up Outputs.The underlined bold values are the best score.
Model Name Avg. Similarity score
Deepseek-R1 78.7
Deepseek-chat 79.4
Gemini 2.5 Pro Preview (03-25) 86.5
Gemini 2.0 Flash 66.5
Llama 3.3 70B 75.0
QWQ-32 83.7
Gemini 2.5 Flash Preview (04-17, Thinking) 82.6
Gemini 2.5 Flash Preview (04-17) 80.2
We generated model outputs using the Engineering template 3 and the Evaluation template 4 separately. As shown in Table 7, all eight models exhibited varied performance on generating treatment follow-up outputs, with average similarity scores ranging from (66.5 %) to (86.5 %). The Gemini 2.5 Pro Preview (03-25) led the group at (86.5 %), followed by QWQ-32 at (83.7 %) and the Gemini 2.5 Flash Preview (04-17, Thinking) at (82.6 %). The secondary Flash Preview (04-17) achieved (80.2 %), while Deepseek-chat and Deepseek-R1 produced (79.4 %) and (78.7 %), respectively. Llama 3.3 70B score (75.0 %), and the earliest variant, Gemini 2.0 Flash, trailed at (66.5 %). The substantial (20.0 %) spread between the top and bottom performers highlights that follow-up coherence poses a greater challenge than initial treatment or diagnosis tasks. Gemini 2.5 Pro Preview’s strong showing suggests its latest context-tracking optimizations significantly enhance continuity in multi-turn recommendations. QWQ-32’s robust (83.7%) score further underscores its ability to maintain patient context across successive exchanges. By contrast, Gemini 2.0 Flash’s steep drop to (66.5 %) indicates limitations in its reasoning architecture and narrower context window when managing follow-up prompts. The approximately (16.1 %) improvement from Gemini 2.0 Flash to the Gemini 2.5 Flash Preview (04-17, Thinking) illustrates the clear benefit of the "Thinking" inference mode for preserving treatment trajectories. Meanwhile, Deepseek-chat and Deepseek-R1, both below the (80.0% ) mark, demonstrate competence with basic follow-up queries but lack the advanced continuity afforded by newer model iterations. Llama 3.3 70B’s (75.0% ) result similarly suggests that even large foundational architectures may struggle with nuanced multi-turn follow-through without specialized enhancements.
6.4 Classification
Table 8: Classification Results: Disorder categories and specific disorders classification dataset evaluation on the chosen models. The underlined bold values are the best score for the specific column.
LLM Categories Disorder Specific Disorders
F1-Weighted Subset Accuracy F1-Weighted Subset Accuracy
DeepSeek R1 0.72 0.54 0.50 0.44
DeepSeek Chat 0.63 0.31 0.41 0.22
Gemini 2.5 Pro Preview (03-25) 0.60 0.26 0.42 0.23
Gemini 2.0 Flash 0.63 0.31 0.46 0.31
Llama 3.3 70B 0.68 0.46 0.47 0.34
QWQ 32 0.65 0.38 0.45 0.33
Gemini 2.5 Flash (04-17, Thinking) 0.60 0.26 0.37 0.14
Gemini 2.5 Flash (04-17) 0.68 0.47 0.44 0.36
We used the Engineering prompt template 5 to generate the model outputs, and then evaluated them for accuracy in classification without using a separate evaluation prompt.As shown in Table 8 compares each model’s ability to assign both broad disorder categories and granular specific disorders using two metrics: F1-weighted score and subset accuracy. Across the board, models achieve higher scores on the broader category task F1-weighted scores range from (0.60) to (0.72), while subset accuracy spans (0.26) to (0.54). In contrast, performance on specific disorders is uniformly lower, with F1-weighted values between (0.37) and (0.50) and subset accuracy from (0.14) to (0.44). This pattern reflects the greater difficulty of exactly matching the full set of fine-grained labels for each case.
DeepSeek-R1 leads in the category classification task, posting the highest F1-weighted score of (0.72) and subset accuracy of (0.54). Close behind, Llama 3.3 70B and the Gemini 2.5 Flash (04-17) variant both achieve balanced category performance, Llama at (0.68)/(0.46) and Gemini Flash at (0.68)/(0.47). For specific disorders, DeepSeek-R1 again tops the field with (0.50) F1-weighted and (0.44) subset accuracy, indicating its relative strength in identifying multiple, co-occurring disorders. QWQ-32 and Gemini 2.0 Flash also perform respectably on the specific task, with F1-weighted scores of (0.45)-(0.46).
The gap between category and specific-disorder scores, averaging around (0.20)-(0.25) F1 points, highlights the challenge of multi-label, fine-grained classification in clinical contexts. Models that excel at broad categorization still struggle to capture every nuanced diagnosis in a single pass, as evidenced by the lower subset accuracies across the board. These results underscore the need for targeted improvements, whether through additional domain data, hierarchical label modeling, or specialized prompt strategies to push LLMs toward clinically reliable, multi-label diagnostic performance.
6.5 Management Plan
Table 9: Management Plan Results: Average similarity scores of model-generated clinical management plans. The underlined bold values are the best score.
Model Name Avg. Similarity score
Deepseek-R1 84.4
Deepseek-chat 83.5
Gemini 2.5 Pro Preview (03-25) 84.1
Gemini 2.0 Flash 78.7
Llama 3.3 70B 82.1
QWQ-32 82.9
Gemini 2.5 Flash Preview (04-17, Thinking) 82.4
Gemini 2.5 Flash Preview (04-17) 81.1
We generate model outputs using the Engineering prompt template 6 and the Evaluation prompt template 7 separately.As shown in Table 9, all eight models produced high-quality management plans, with average similarity scores ranging from (78.7 %) to (84.4 %). This narrow spread of (5.7 %) points indicates that even the lowest-scoring model generates outputs that align reasonably well with expert-authored references, while the top performers exhibit only modest gains over the rest.DeepSeek-R1 achieved the highest score at (84.4%), marginally outperforming Gemini 2.5 Pro Preview (03-25) at (84.1 %) and DeepSeek-chat at (83.5 %). These leading models consistently synthesized comprehensive therapeutic recommendations, patient education points, and follow-up strategies that closely matched the reference plans. The strong performance of Gemini 2.5 Pro Preview underscores the impact of its recent architectural refinements on complex planning tasks. Mid-tier models QWQ-32 at (82.9 %), Gemini 2.5 Flash Preview (04-17, Thinking) at (82.4 %), and Llama 3.3 70B at (82.1 %) also deliver robust outputs, suggesting that both reasoning enhancements ("Thinking" mode) and broad contextual training contribute to solid plan generation. Gemini 2.5 Flash Preview (04-17) score (81.1 %), while Gemini 2.0 Flash lagged at (78.7 %), highlighting the clear benefit of iterative updates and instruction-tuning in newer model versions. Despite high similarity scores overall, careful human review remains essential to catch any subtle omissions or safety concerns before clinical use.
6.6 Clinical Approach
We generated model outputs using the Engineering prompt template 8 and the Evaluation prompt template 9 separately. As shown in Table 10 summarizes each model’s performance on the Clinical Approach task under two prompt formulations. Under the original prompt (Approach 1), average similarity scores range from (86.7 %) to (90.1 %), demonstrating that all models can reliably outline diagnostic frameworks and management plans when given a clear, familiar instruction. With the reformulated prompt (Approach 2), scores fall between (82.8 %) and (87.5 %), indicating a consistent sensitivity to prompt wording across architectures. Deepseek-R1 and Deepseek-chat start at the top of the leaderboard with (90.0 %) and (90.1 %) under Approach 1, respectively, but both experience moderate declines of (4.1 %) and (4.7 %) when switched to Approach 2. Llama 3.3 70B drops from (89.7 %) to (83.9 %) (- 5.8 %), while QWQ-32 falls from (89.2 %) to (82.8 %) (- 6.4 %), marking the largest sensitivity among the group. These results suggest that even high-capacity models can struggle to generalize reasoning steps if the prompt structure shifts.By contrast, Gemini 2.5 Pro Preview (03-25) exhibits the greatest robustness, declining only from (89.2 %) to (87.5 %) (- 1.7 %). Its Flash variants also remain relatively stable, with the "Thinking" mode decreasing by (2.1 %) (87.5 % → 85.4 %) and the standard Flash by (3.5 %) (88.6 % → 85.1 %). This relative consistency underscores the value of instructional tuning and specialized inference strategies in maintaining performance when prompt formulations vary.
Table 10: Clinical Approach Results: Average similarity scores for different language models across two prompt formulations on the Clinical Approach task. The "Change" column indicates the difference between Approach 2 and Approach 1. The underlined bold values are the best score for the specific column.
Model Name Avg. Similarity Score
Approach 1 Approach 2
Deepseek-R1 90.0 85.9
Deepseek-chat 90.1 85.4
Gemini 2.5 Pro Preview (03-25) 89.2 87.5
Gemini 2.0 Flash 86.7 85.1
Llama 3.3 70B 89.7 83.9
QWQ-32 89.2 82.8
Gemini 2.5 Flash Preview (04-17, Thinking) 87.5 85.4
Gemini 2.5 Flash Preview (04-17) 88.6 85.1
6.7 Mental QA
Table 11: Mental QA Results: Average similarity scores for different language models across two prompt versions on the Mental QA task. The "Change" column represents the difference between Prompt 1 and Prompt 2. The underlined bold values are the best score for the specific column.
Model Name Avg. Similarity Score
Prompt 1 Prompt 2
Deepseek-R1 88.0 88.1
Deepseek-chat 82.7 88.1
Gemini 2.5 Pro Preview (03-25) 87.0 84.6
Gemini 2.0 Flash 82.7 84.0
Llama 3.3 70B 84.2 84.1
QWQ-32 86.0 84.3
Gemini 2.5 Flash Preview (04-17, Thinking) 79.3 85.9
Gemini 2.5 Flash Preview (04-17) 83.5 83.1
We use the Engineering prompt template 10. As shown in Table 11, models’ average similarity scores under Prompt 1 range from (79.3 %) to (88.0 %), while under Prompt 2 they span (83.1 %) to (88.1 %). Deepseek-R1 remains near the top of both distributions with (88.0 %) and (88.1 %), respectively, and Deepseek-chat jumps from (82.7 %) to (88.1 %). At the lower end, the Gemini 2.5 Flash Preview (04-17, Thinking) rises from (79.3 %) to (85.9 %), whereas the Gemini 2.5 Pro Preview (03-25) declines from (87.0 %) to (84.6 %). These shifts underscore that definitional accuracy and its alignment with expert references varies substantially with prompt formulation.The magnitude of change varies by architecture. Deepseek-chat’s sizable +5.4 point gain and the Gemini 2.5 Flash (Thinking) model’s +6.6 point improvement highlight how enhanced instructional phrasing can bolster performance even in fact-based tasks. Conversely, the Gemini 2.5 Pro Preview’s -2.4 point drop demonstrates that a seemingly minor prompt tweak can grace alignment for some models. QWQ-32 (86.0 % → 84.3 %) and LLaMA 3.3 70B (84.2 % → 84.1 %) show minimal (<1 point) fluctuations, indicating greater prompt robustness in these cases.Deepseek-R1’s near-identical scores across both prompts suggest it generalizes definitional reasoning most consistently. Overall, these results emphasize the critical role of prompt Engineeringing in terminology-focused QA: small changes in phrasing or structure can yield double-digit performance swings. For sensitive domains like psychiatry, ensuring prompt clarity and close alignment with a model’s instruction-following tendencies is essential to achieve reliable, clinically accurate definitions.
6.8 Sequential Question Answering
Table 12: Sequential Question Answering Results: Average similarity scores for various language models on the Sequential Question Answering task using the baseline prompt (Prompt 1). The underlined bold values are the best score.
Model Name Avg. Similarity Score
Deepseek-R1 89.0
Deepseek-chat 88.9
Gemini 2.5 Pro Preview (03-25) 96.9
Gemini 2.0 Flash 90.0
Llama 3.3 70B 81.4
QWQ-32 86.2
Gemini 2.5 Flash Preview (04-17, Thinking) 89.7
Gemini 2.5 Flash Preview (04-17) 89.2
We generated model outputs using the Engineering prompt template 11 and the Evaluation prompt template12 separately.As shown in Table 12, all eight models exhibit strong performance on the Sequential Question Answering task under Prompt 1, with average similarity scores ranging from (81.4 %) to (96.9 %). Gemini 2.5 Pro Preview (03-25) stands out with an exceptional (96.9 %), demonstrating its superior ability to maintain context and deliver coherent, expert-aligned answers across multiple turns. Deepseek-R1 and Deepseek-chat follow closely at (89.0 %) and (88.9 %), respectively, indicating robust sequential reasoning capabilities.Mid-tier models including Gemini 2.5 Flash Preview (04-17, Thinking) at (89.7 %), Gemini 2.5 Flash Preview (04-17) at (89.2 %), and Gemini 2.0 Flash at (90.0 %) also deliver consistently high-quality outputs, showcasing effective context tracking and answer continuity. QWQ-32 achieves a solid (86.2 %), reinforcing its reliability in multi-turn clinical dialogues. These scores suggest that both "Thinking" inference modes and incremental flash updates materially enhance a model’s sequential QA performance. Llama 3.3 70B records the lowest similarity at (81.4 %), highlighting potential difficulties in capturing all nuanced dependencies between successive questions without additional prompt optimization. Nevertheless, its performance remains respectable, and prior analyses suggest that refined prompt Engineeringing could further close the gap. Overall, these results underscore that while top-tier proprietary models excel in sequential reasoning, prompt clarity and model updates play a critical role in elevating performance across the board.
6.9 Multiple Choice Questions (MCQ)
This section presents model performance on structured, objective assessment items commonly used in medical education. We start with single-best-answer questions (MCQs), followed by more complex matching and exam-style prompts. These formats test discrete factual recall, application of clinical reasoning, and the ability to integrate information across multiple elements.
6.9.1 Single-Answer Multiple-Choice Question
We used the Engineering prompt template 13 to generate the model outputs, and then evaluated them for accuracy in classification without using a separate evaluation prompt.As shown in Table 13, model performance on the MCQ task varies considerably depending on prompt style. Under the baseline prompt (P1), accuracy ranges from (75.5 %) to (83.2 %) and balanced accuracy from (64.3 %) to (79.3 %). With the few-shot prompt (P2), overall accuracy spans (71.6 %) to (82.8 %) and balanced accuracy (63.7 %) to (74.4 %). This indicates that while some models maintain consistency across both formats, others particularly those sensitive to prompt examples experience notable drops.Gemini 2.5 Pro Preview (03-25) delivers the most robust performance overall, scoring (81.7 %)/(79.3 %) in P1 accuracy/BA and improving to (82.8 %)/(74.4 %) in P2. Deepseek-R1 achieves the highest P1 accuracy at (83.2 %) but falls to (71.6 %) under P2, suggesting a strong grasp of standalone questions but shows better generalization when provided with examples. Llama 3.3 70B and QWQ-32 show stable results around (77.1 %)/(72.2 %) and (78.4 %)/(74.8 %) in P1, and (77.5 %)/(72.3 %) and (76.8 %)/(71.5 %) in P2 highlighting reliable reasoning unaffected by prompt format.By contrast, Gemini 2.0 Flash underperforms relative to its successors, with balanced accuracy at (64.3 %)/(65.2 %) in P1/P2 despite moderate accuracy (75.59 %)/(75.2 %). The Gemini 2.5 Flash Preview variants also show mixed results: the "Thinking" mode drops to (74.1 %)/(63.7 %) in P2, while the standard Flash holds (79.1 %)/(71.5 %). These findings underscore that MCQs though simpler than open-ended tasks still reveal differences in how models leverage prompt contexts and example-based cues to select correct answers.
Table 13: Multiple-choice question Results: Accuracy (A) and balanced accuracy (BA) scores for each model across all books and prompts.The underlined bold values are the best score for the specific column.
Model Name P1 P2
A BA A BA
Deepseek-R1 83.2 76.4 71.6 72.3
Deepseek-chat 75.8 71.5 74.4 69.1
Gemini 2.5 Pro Preview (03-25) 81.7 79.3 82.8 74.4
Gemini 2.0 Flash 75.5 64.3 75.2 65.2
Llama 3.3 70B 77.0 72.2 77.5 72.3
QWQ-32 78.4 74.8 76.8 71.5
Gemini 2.5 Flash Preview (04-17, Thinking) 79.4 69.8 74.1 63.75
Gemini 2.5 Flash Preview (04-17) 79.5 69.8 79.1 71.5
6.9.2 Extended Matching Item
[…]

### v2

@@ 5 Results & Discussion
This section presents a multi-faceted analysis of the performance of 15 leading LLMs on the PsychiatryBench benchmark. Our evaluation, summarized in Table 6, provides a granular view of each model’s capabilities across eleven distinct diagnostic reasoning tasks. Complementing this, Figure 3 offers a macroscopic perspective, plotting each model’s average performance against its release date, with bubble size representing the model’s parameter count. This integrated approach allows for a thorough discussion of overarching performance trends, the impact of model scale, the value of domain-specific training, and the efficacy of different inference strategies.
5.1 Performance of Medical Models
In this section, we evaluate the medical models introduced in Section 4.1.2 to establish a clinically meaningful baseline. Table 5 reports the performance of several specialized medical LLMs alongside multiple frontier models. II-Medical-8B attains (76.8%) on both Management Plan and Mental QA, indicating well-balanced performance across tasks. OpenBioLLM achieves (75.1%) on Management Plan and a higher (82.2%) on Mental QA. Med-Palmyra obtains similarly competitive results, with (78.2%) and (83.2%), respectively. In contrast, JSL-MedLlama performs substantially worse, scoring (61.0%) on Management Plan and (65%) on Mental QA, falling well below the 75% threshold and therefore being excluded from subsequent specialized analyses. Med-PaLM 2 was also excluded due to deployment constraints that require formal Google approval for medical use cases, which prevented local fine-tuning and evaluation.
The underlined bold entries in Table 5 denote the best-performing medical models within this subset, rather than the best overall scores. This distinction is important because several general-purpose frontier LLMs evaluated elsewhere in the manuscript achieve markedly higher performance on PsychiatryBench, surpassing even the strongest medical models. Accordingly, the underline identifies the top results within the medical-model category only, not the global optimum across all models considered.
We further evaluated the medical language models using two complementary benchmark datasets, Management Plan and Mental QA, each targeting distinct aspects of clinical competence. Together, these datasets yield a balanced assessment of practical clinical decision-making and domain-specific psychiatric knowledge, providing a comprehensive characterization of medical LLM performance.
Table 5: Evaluation of medical and frontier LLMs on clinical diagnosis and follow-up tasks using benchmark datasets. The underlined bold values indicate the best score specifically among the listed medical models.
Models Management Plan Mental QA
II-Medical-8B 76.8 76.8
OpenBioLLM 75.1 82.2
MedGemma 81.7 87.1
Med-Palmyra 78.2 83.2
JSL-MedLlama 61.0 65.0
Gemini 2.5 Flash 81.1 83.5
Gemini 2.5 Flash (T) 82.4 79.3
Deepseek-chat 83.5 82.7
Among all medical LLMs examined, MedGemma attains the strongest overall results, with (81.7%) on Management Plan and (87.1%) on Mental QA. These clinical benchmarks also help contextualize the notably stronger performance of general-purpose frontier LLMs. For example, Deepseek-chat reaches (83.5%) on Management Plan, and Gemini 2.5 Flash (T) achieves (82.4%), both outperforming MedGemma. While the medical models exhibit strong biomedical knowledge and structured reasoning, the frontier LLMs demonstrate superior coherence, adaptability, and open-ended reasoning capabilities that are particularly valuable for psychiatric assessment. In our benchmarking, we chose MedGemma to be representative to medical models in the benchmarking against other SoTA LLMs.
5.2 Overall Performance Landscape and a Clear Trajectory of Improvement
The results unequivocally demonstrate a stratified performance landscape and a rapid, positive trajectory of advancement in LLM capabilities for diagnostic reasoning. As detailed in Table 6, a distinct top tier of models has emerged. GPT 5 Medium (T) stands out as the premier model with the highest average score of (84.5%), closely followed by Sonnet 4.5 (T) (83.7%). These models consistently deliver state-of-the-art performance, securing the best or second-best scores across the majority of tasks, particularly those involving complex clinical judgment like Diagnosis, Treatment, and management plan.
Conversely, a clear performance gap separates these leaders from earlier and smaller models. Gemini 2.0 Flash (74.6%) and GPT-oss (72.3%) anchor the lower end of the performance spectrum. Figure 3 visually articulates this evolutionary trend, revealing a strong positive correlation between model release date and average performance. For instance, average scores improve by over +8 points between Gemini 2.0 Flash (74.6%) and the Sonnet 4.5 series (83.5% –83.7%), and by more than +10 points relative to baselines such as GPT-oss (72.3%). The jump of +4.0 points from Sonnet 4 (80.5%) to GPT 5 Medium (T) (84.5%) in a matter of months underscores the accelerated pace of innovation. This progression suggests that continuous refinements in model architecture, training methodologies, and data curation are yielding substantial and measurable improvements in clinical applicability.
5.3 Task-Specific Performance Analysis: Identifying Strengths and Persistent Challenges
A horizontal analysis of Table 6 reveals consistent patterns in task difficulty across models. Modern LLMs demonstrate remarkable proficiency in tasks that require synthesizing contextual information and generating structured, long-form clinical reasoning. This is most evident in Sequential QA, where Sonnet 4.5 (T) achieved a near-perfect score of (96.2%). Similarly, strong performance in the Clinical Approach task topped at (90.2%) by Sonnet 4.5 shows that contemporary models are adept at constructing coherent diagnostic and management pathways from complex psychiatric vignettes.
Conversely, two task categories remain persistent challenges, exposing the limits of current LLM capabilities. First, the Classification of Specific Disorders task proved to be the most difficult across the benchmark, with even top-tier models like GPT 5 Medium (T) achieving only (0.52/45.0) F1-score/subset-accuracy. This reflects the inherent difficulty of multi-label classification in psychiatry, where overlapping symptoms and comorbidities blur categorical boundaries. Second, the Extended Matching Items (EMI) task, which demands discriminating among numerous clinically similar options, also showed broad performance variability. While models like GPT 5 Medium (T) scored (89.1%), Gemini 2.0 Flash recorded a much lower score of (75.5%) in this task. Separately, on the Exams task, Gemini 2.0 Flash had one of the lowest results at (69.9%), underscoring broad difficulties in structured assessment tasks. These findings highlight that while LLMs have mastered structured narrative reasoning, fine-grained clinical discrimination remains a frontier for further optimization.
5.4 Specialized vs. Generalist Models: The Case of MedGemma
A compelling narrative within our results is the performance of the domain-specialized model, MedGemma. Despite its moderate size, MedGemma achieves an impressive average score of (78.5%), placing it on par with large-scale generalist models like Gemini 2.5 Pro (80.2%) and DeepSeek-R1 (80.4%). This performance is not uniform; rather, it is concentrated in areas that directly benefit from its specialized training on biomedical and clinical texts.
As highlighted in Table 6, MedGemma’s distinct advantage is evident in its strong F1/Subset Accuracy scores in the highly granular Classification of Specific Disorders task (0.69/45.0), where it outperforms nearly all generalist counterparts. Furthermore, it achieved one of the highest results in the knowledge-intensive MCQ task (87.4%), reflecting superior factual retention and domain recall.
However, this specialization comes with a trade-off, as MedGemma was less competitive in broader, open-ended reasoning tasks such as Management Plan (81.7%) and Sequential QA (87.7%) compared to top-tier generalist models like Sonnet 4.5 (T) (96.2%) and GPT 5 Medium (T) (96.1%). This suggests a dichotomy where generalist models excel at fluid, multi-step reasoning and contextual integration, while specialized models remain superior in precision-driven, knowledge-based clinical classification and recall.
Table 6: Performance comparison of 15 selected LLMs across all PsychiatryBench tasks. Classification tasks are evaluated using F1 Score and Subset Accuracy (F1/A), while other tasks are assessed based on the similarity between model responses and reference answers, as judged by an LLM evaluator. (T) denotes Thinking models. Green, blue, and red highlight the best, second-best, and lowest scores per column, respectively.
Model Diagnosis Treatment Treatment Follow-Up Classification Management Plan Clinical Approach Mental QA Sequential QA MCQ Exams EMI AVG.
Categories Specific
DeepSeek -R1
84.2 79.9 78.7 0.72/54.0 0.50/44.0 84.4 90.0 88.0 89.0 83.2 79.2 85.8 80.4
DeepSeek -chat
82.7 81.9 79.4 0.63/31.0 0.41/23.0 83.5 90.1 82.7 88.9 75.8 71.4 79.8 76.7
Gemini 2.0 Flash
83.1 77.4 66.5 0.63/31.0 0.46/31.0 78.7 86.7 82.7 90.0 75.6 69.9 75.5 74.6
Gemini 2.5 Pro
86.0 80.5 86.5 0.60/26.0 0.42/31.0 84.1 89.2 87.0 96.9 81.7 80.4 87.9 80.2
Gemini 2.5 Flash
78.0 81.0 80.2 0.68/47.0 0.44/36.0 81.1 88.6 83.5 89.2 79.5 73.1 84.5 77.6
Gemini 2.5 Flash (T)
85.0 82.9 82.6 0.60/26.0 0.37/14.0 82.4 87.5 79.3 89.7 79.4 78.4 86.1 77.5
Llama 3.3 70B
83.6 82.3 75.0 0.68/46.0 0.47/34.0 82.1 89.7 84.2 81.4 77.1 66.5 72.2 75.8
QWQ-32 85.9 82.1 83.7 0.65/38.0 0.45/33.0 82.9 89.2 86.0 86.2 78.4 73.4 79.8 78.1
QWEN 3 32B
82.1 81.7 80.9 0.58/25.0 0.35/14.0 80.3 90.0 90.6 89.7 74.2 65.3 73.1 75.1
MedGemma 79.4 84.2 79.2 0.55/34.0 0.69/45.0 81.7 90.0 87.1 87.7 87.4 69.2 72.5 78.5
GPT-oss 70.8 67.8 64.5 0.68/53.0 0.46/39.0 72.2 73.5 86.3 93.4 74.6 72.2 78.4 72.3
GPT 5 Medium (T) 88.2 88.0 90.0 0.75/59.0 0.52/45.0 87.9 90.1 90.7 96.1 87.1 79.8 89.1 84.5
Sonnet 4 85.8 84.5 83.5 0.69/45.0 0.47/34.0 84.8 90.0 89.4 89.6 81.6 77.0 83.9 80.5
Sonnet 4.5 89.6 86.6 89.3 0.70/43.0 0.44/30.0 87.0 90.2 81.8 95.6 83.0 75.6 84.9 81.5
Sonnet 4.5 (T)
89.9 88.4 90.4 0.73/55.0 0.40/33.0 88.6 87.5 92.5 96.2 89.2 80.0 88.9 83.7
Figure 3: Each bubble shows a model, positioned by release date and average performance. Bubble size denotes parameter count, and color represents model family. Larger, newer models generally achieve higher performance.
5.5 The Role of Inference Strategies: "Thinking" vs. Standard Modes
Our study’s inclusion of "Thinking" variants for the Gemini 2.5 Flash and Sonnet 4.5 models reveals that the benefit of more deliberative, multi-step inference is highly architecture-dependent. For the Anthropic models, this strategy yielded a significant performance dividend. Sonnet 4.5 (T) not only surpassed its standard counterpart with an average score of (83.7%) versus (81.5%) but also achieved the highest performance across several of the most cognitively demanding tasks: Diagnosis (89.9%), Treatment (88.4%), Treatment Follow-Up (90.4%), and Mental QA (92.5%). This indicates that for the Sonnet architecture, the “Thinking” mode effectively activates deeper and more structured clinical reasoning pathways.
In contrast, this advantage was not evident for the Google Gemini models. The average performance of Gemini 2.5 Flash (T) (77.5%) was negligibly lower than that of its standard counterpart, Gemini 2.5 Flash (77.6%). Moreover, the ’Thinking’ variant’s performance was mixed: it underperformed in Mental QA (79.3%) compared to the standard model’s (83.5%), but it scored higher in Diagnosis (85.0%) versus the standard model’s (78.0%), indicating an inconsistent benefit from the additional inference step.
5.6 The Impact of Model Scale and Release Date
The scatter plot in Figure 3 provides critical insights into the relationship between model size (parameter count), recency, and overall performance. A clear upward trend emerges: larger models depicted by larger bubbles tend to cluster in the higher performance range. The leading results achieved by Sonnet 4.5 (T) and GPT 5 Medium (T), with average scores of (83.7%) and (84.5%) respectively, exemplify how scale combined with architectural refinement drives superior clinical reasoning outcomes.
However, size alone does not guarantee success. The performance of DeepSeek-R1, a large-scale model with an average score of (80.4%), remains strong yet is surpassed by more recent and structurally optimized models like Sonnet 4.5 (T) and GPT 5 Medium (T). This pattern underscores that architectural innovation, training corpus diversity, and post-alignment strategies can yield greater performance gains than mere parameter expansion.
The overall trajectory visible in Figure 3 confirms this trend: newer architectures are consistently leveraging scale more effectively, converting size into meaningful improvements in reasoning accuracy, clinical adaptability, and task generalization.
5.7 Model Family Performance
Our analysis reveals distinct performance signatures across model families, reflecting different architectural priorities and the effectiveness of their training. The frontier generalist models from Anthropic and OpenAI set the benchmark, with Sonnet 4.5 (T) (83.7%) and GPT-5 Medium (T) (84.5%) demonstrating state-of-the-art capabilities driven by recent architectural innovations.
In contrast, other families highlight key limitations. The Gemini models, while strong with Gemini 2.5 Pro at 80.2%, show that advanced inference modes are not universally beneficial; its ’Thinking’ variants yielded inconsistent gains, unlike the significant boosts seen in the Sonnet series. Similarly, DeepSeek-R1 (80.4%) illustrates that massive scale alone does not guarantee leadership, as it is surpassed by more structurally optimized models.
Finally, specialized medical models like MedGemma (78.5%) confirm a classic trade-off: their domain-specific pretraining provides a measurable edge in knowledge-intensive tasks but limits the cognitive flexibility needed for nuanced psychiatric reasoning. These patterns confirm that the frontier of performance is now defined more by sophisticated architectural design and alignment than by raw parameter count alone.
5.8 Cross-Task Consistency and Stability Analysis
[…]